calculate the mass of potassium chlorate required to librate 6.72 litere of oxygen at STP
1
Expert's answer
2019-08-23T04:18:53-0400

6.72 liters of oxygen is:

"\frac {6.72L}{22.4 mol/L} = 0.3 mol" , here we assume that O behaves as an ideal gas and we take 1 mole of gas as 22.4mol/L.


By the reaction below molar ratio O2:KClO3 is 3:2

"2KClO_3 = 2KCl + 3O_2"

So, by stoichiometry of the reaction we can find moles of potassium chlorate:

"0.3 moles \cdot \frac {2moles}{3moles}=0.2 moles"

mass of potassium chlorate is:

"0.2 moles \cdot 122.5 g/mol = 24.5 g"